OpenStudy (anonymous):

How do I even start this question: Grace is three years younger than her sister Jane. Twice the sum of their ages is 4 more than their father's age. if their father is 54, find the ages of Grace and Jane.

OpenStudy (anonymous):

\[g=j-3\] \[2(g+j)=4+father\]

OpenStudy (anonymous):

When it says 4 more, does that mean 4 times or just adding on 4?

OpenStudy (anonymous):

i have four more candy's than you???

OpenStudy (anonymous):

it means +

OpenStudy (anonymous):

if it is times it will say times or product

OpenStudy (anonymous):

Oh ok

OpenStudy (anonymous):

So after doing that, what do I need to do?

OpenStudy (anonymous):

solve simultaniously \[g=3+j\] \[2(g+j)=4+54=58\] substitute g in to equation (2)

OpenStudy (anonymous):

oh ok